OpenStudy (anonymous):

A rectangular shed covers 800 square feet of ground. It is 10 feet longer than it is wide. What is the width of the shed to the nearest tenth of a foot?

OpenStudy (jhannybean):



OpenStudy (jhannybean):

So \(A =lw\)\[\begin{align} A = w \cdot (10+w) \\&: ~A = 10w +w^2 \\&: ~800 = 10w+w^2 \\&: ~w^2+10w-800=0 \end{align}\]First check to see if the discriminant of this quadratic is positive. If it is, then you can state that you have 2 solutions. \[\begin{align} b^2-4ac = (10)^2 -4(1)(-800) = 3300 \end{align}\]This means you have 2 solutions.

OpenStudy (perl):

one of the solutions is negative, so you can reject it

OpenStudy (jhannybean):

\[\begin{align}\ w^2+10w-800=0 \\&: (w^2+10w)-800=0 \\&: (w^2+10w+25\color{red}{-25})-800\color{red}{-25}=0 \\&: (w+5)^2 -825=0 \\&: (w+5)^2 = 825 \\&: w+5 = \pm\sqrt{825} \\&: w = -5 \pm \sqrt{825} \end{align}\]

OpenStudy (jhannybean):

So tell me which value of w is positive :)